# GMAI-MMBench: A Comprehensive Multimodal Evaluation Benchmark Towards General Medical AI

### v4

@@ E.2 Case Study
| 10 | MR | H | Image Level | Correct |
| 11 | C | H | Image Level | Correct |
| 12 | SWR | ENT | Image Level | Correct |
| 13 | DD | GH | Image Level | Correct |
| 14 | ASR | NH | Image Level | Correct |
| 15 | SAR | U | Box Level | Correct |
| 16 | DD | PM | Box Level | Correct |
| 17 | OR-NH | E | Mask Level | Correct |
| 18 | OR-P | U | Contour Level | Correct |
| 19 | SIR | GS | Box Level | Correct |
| 20 | BVR | H | Mask Level | Correct |
| 21 | CR | H | Box Level | Correct |
| 22 | DD | CS | Mask Level | Correct |
| 23 | DD | OS | Contour Level | Correct |
| 24 | NT | O | Mask Level | Correct |
| 25 | OR-T | PM | Mask Level | Correct |
| 26 | SIR | GS | Mask Level | Correct |
| 27 | SWR | GS | Image Level | Question misunderstanding |
| 28 | BVR | O | Mask Level | Question misunderstanding |
| 29 | ACR | OS | Mask Level | Question misunderstanding |
| 30 | MR | GH | Image Level | Question misunderstanding |
| 31 | C | H | Image Level | Perceptual error |
| 32 | SWR | GS | Image Level | Perceptual error |
| 33 | OR-T | PM | Mask Level | Perceptual error |
| 34 | AR | LMP | Image Level | Perceptual error |
| 35 | NT | N | Mask Level | Perceptual error |
| 36 | DD | CS | Box Level | Perceptual error |
| 37 | DD | D | Mask Level | Perceptual error |
| 38 | DD | GH | Contour Level | Perceptual error |
| 39 | OR-T | PM | Mask Level | Perceptual error |
| 40 | NT | N | Mask Level | Perceptual error |
| 41 | OR-T | PM | Contour Level | Perceptual error |
| 42 | DD | O | Image Level | Lack of knowledge |
| 43 | IQG | O | Image Level | Lack of knowledge |
| 44 | MR | LMP | Image Level | Lack of knowledge |
| 45 | SAR | GS | Box Level | Lack of knowledge |
| 46 | SAR | U | Box Level | Lack of knowledge |
| 47 | DD | PM | Mask Level | Lack of knowledge |
| 48 | NT | O | Mask Level | Lack of knowledge |
| 49 | SG | LMP | Image Level | Lack of knowledge |
| 50 | DD | O | Image Level | Lack of knowledge |
| 51 | SG | LMP | Image Level | Lack of knowledge |
| 52 | DD | OM | Image Level | Lack of knowledge |
| 53 | AR | GS | Image Level | Lack of knowledge |
| 54 | AR | OG | Image Level | Lack of knowledge |
| 55 | DD | D | Image Level | Lack of knowledge |
| 56 | DD | U | Image Level | Lack of knowledge |
| 57 | DD | OS | Image Level | Irrelevant response |
| 58 | AR | ID | Image Level | Irrelevant response |
| 59 | AR | OS | Image Level | Irrelevant response |
| 60 | ASR | OG | Image Level | Irrelevant response |
| 61 | DD | PM | Image Level | Refuse to answer |
| 62 | BVR | O | Mask Level | Refuse to answer |

### v6

@@ E.2 Case Study
10 | MR | H | Image Level | Correct |
11 | C | H | Image Level | Correct |
12 | SWR | ENT | Image Level | Correct |
13 | DD | GH | Image Level | Correct |
14 | ASR | NH | Image Level | Correct |
15 | SAR | U | Box Level | Correct |
16 | DD | PM | Box Level | Correct |
17 | OR-NH | E | Mask Level | Correct |
18 | OR-P | U | Contour Level | Correct |
19 | SIR | GS | Box Level | Correct |
20 | BVR | H | Mask Level | Correct |
21 | CR | H | Box Level | Correct |
22 | DD | CS | Mask Level | Correct |
23 | DD | OS | Contour Level | Correct |
24 | NT | O | Mask Level | Correct |
25 | OR-T | PM | Mask Level | Correct |
26 | SIR | GS | Mask Level | Correct |
27 | SWR | GS | Image Level | Question misunderstanding |
28 | BVR | O | Mask Level | Question misunderstanding |
29 | ACR | OS | Mask Level | Question misunderstanding |
30 | MR | GH | Image Level | Question misunderstanding |
31 | C | H | Image Level | Perceptual error |
32 | SWR | GS | Image Level | Perceptual error |
33 | OR-T | PM | Mask Level | Perceptual error |
34 | AR | LMP | Image Level | Perceptual error |
35 | NT | N | Mask Level | Perceptual error |
36 | DD | CS | Box Level | Perceptual error |
37 | DD | D | Mask Level | Perceptual error |
38 | DD | GH | Contour Level | Perceptual error |
39 | OR-T | PM | Mask Level | Perceptual error |
40 | NT | N | Mask Level | Perceptual error |
41 | OR-T | PM | Contour Level | Perceptual error |
42 | DD | O | Image Level | Lack of knowledge |
43 | IQG | O | Image Level | Lack of knowledge |
44 | MR | LMP | Image Level | Lack of knowledge |
45 | SAR | GS | Box Level | Lack of knowledge |
46 | SAR | U | Box Level | Lack of knowledge |
47 | DD | PM | Mask Level | Lack of knowledge |
48 | NT | O | Mask Level | Lack of knowledge |
49 | SG | LMP | Image Level | Lack of knowledge |
50 | DD | O | Image Level | Lack of knowledge |
51 | SG | LMP | Image Level | Lack of knowledge |
52 | DD | OM | Image Level | Lack of knowledge |
53 | AR | GS | Image Level | Lack of knowledge |
54 | AR | OG | Image Level | Lack of knowledge |
55 | DD | D | Image Level | Lack of knowledge |
56 | DD | U | Image Level | Lack of knowledge |
57 | DD | OS | Image Level | Irrelevant response |
58 | AR | ID | Image Level | Irrelevant response |
59 | AR | OS | Image Level | Irrelevant response |
60 | ASR | OG | Image Level | Irrelevant response |
61 | DD | PM | Image Level | Refuse to answer |
62 | BVR | O | Mask Level | Refuse to answer |